OpenStudy (anonymous):

Given f(x) = x-5 / 3 solve for f^-1(3).

OpenStudy (anonymous):

\[\frac{ x-5 }{ 3 }\]

OpenStudy (anonymous):

\[f ^{-1}(3)\]

ganeshie8 (ganeshie8):

you want to find out for what value of x, the function spits out `3`

ganeshie8 (ganeshie8):

\[\large 3 = \dfrac{x-5}{3}\] solve \(x\)

OpenStudy (anonymous):

14

OpenStudy (anonymous):

hey, you need to find the inverse function first, and only then plug in 3 for x.

ganeshie8 (ganeshie8):

14 is right !

OpenStudy (anonymous):

Ok thank you, i think I just overthink all of this. seems much easier when you explain

OpenStudy (anonymous):

last one, new thread or post here?

ganeshie8 (ganeshie8):

post it here if it is related to this question

OpenStudy (anonymous):

f(x) = (x-5)/ 3 f(x) = (1/3)x-5/3 y = (1/3)x-5/3 x= (1/3)y-5/3 3x= y-5 3x+5= y f^(-1) (x) = 3x+5 f^(-1) (3) = 3(3)+5 =9+5=14 yes 14 is right, but I just didn't see the process for finding the inverse function.

ganeshie8 (ganeshie8):

one question per post is good though :)

OpenStudy (anonymous):

I see what yu did.

OpenStudy (anonymous):

But I mean if the tack was to find f^(-1)(x) first, you wuldn't get away with plugging 3 for y.

OpenStudy (anonymous):

oops, *task

OpenStudy (anonymous):

i'll post a new thread @((*-*)) good job :D

ganeshie8 (ganeshie8):

the task was to find the inverse of 3 so there is no need to find the inverse of entire function *strictly speaking* :P

OpenStudy (anonymous):

Yes, I know. I was just proposing a general approach, although your's is surely much easier.

ganeshie8 (ganeshie8):

:)